### v4

# TUNeS: A Temporal U-Net with Self-Attention for Video-based Surgical Phase Recognition

Abstract
Objective: To enable context-aware computer assistance in the operating room of the future, cognitive systems need to understand automatically which surgical phase is being performed by the medical team. The primary source of information for surgical phase recognition is typically video, which presents two challenges: extracting meaningful features from the video stream and effectively modeling temporal information in the sequence of visual features. Methods: For temporal modeling, attention mechanisms have gained popularity due to their ability to capture long-range dependencies. In this paper, we explore design choices for attention in existing temporal models for surgical phase recognition and propose a novel approach that uses attention more effectively and does not require hand-crafted constraints: TUNeS, an efficient and simple temporal model that incorporates self-attention at the core of a convolutional U-Net structure. In addition, we propose to train the feature extractor, a standard CNN, together with an LSTM on preferably long video segments, i.e., with long temporal context. Results: In our experiments, almost all temporal models performed better on top of feature extractors that were trained with longer temporal context. On these contextualized features, TUNeS achieves state-of-the-art results on the Cholec80 dataset. Conclusion: This study offers new insights on how to use attention mechanisms to build accurate and efficient temporal models for surgical phase recognition. Significance: Implementing automatic surgical phase recognition is essential to automate the analysis and optimization of surgical workflows and to enable context-aware computer assistance during surgery, thus ultimately improving patient care.
[…]
IEEEexample:BSTcontrol
1 Introduction
[…]
Video-based surgical phase recognition refers to automatically analyzing the video stream that is recorded in the operating room to recognize which surgical phase [1] is being performed. Robust and accurate methods for surgical phase recognition are a prerequisite for computer-assisted surgery, where context-aware assistance functions are to be provided automatically [2]. Here, automatic computer assistance could comprise, for example, providing crucial patient-related information at the right time (intra-operatively) or preparing the surgical report (post-operatively).
Current approaches for surgical phase recognition follow a data-driven approach, meaning that they learn from annotated example videos. Typically, two types of deep-learning models are combined: (i) a visual feature extractor, e.g., a Convolutional Neural Network (CNN) or a Vision Transformer (ViT) [3], and (ii) a temporal model, e.g., a Temporal Convolutional Network (TCN) [4, 5] or a Recurrent Neural Network (RNN). An RNN typically consists of Long Short-Term Memory (LSTM) cells [6] or Gated Recurrent Units (GRUs) [7]. Recent temporal models also integrate attention mechanisms [8, 9].
In general, the feature extractor learns to extract descriptive features from individual video frames. Based on that, the temporal model learns to identify meaningful temporal patterns in the sequence of visual features to identify the surgical phases. For handling long surgical videos, the temporal model should be able to capture long-range temporal relationships.
1.0.1 Visual feature extraction
The question of how to obtain suitable features for phase recognition has not been studied extensively so far. Most methods simply finetune a CNN, typically a ResNet [10], on the task of phase recognition after pre-training on a large-scale dataset. Few studies [11, 12] investigated other architectures such as ViTs or spatio-temporal video models. Others explored training a standard CNN either on additional tasks, e.g., tool recognition [13, 14], surgical step recognition [15], or surgical scene segmentation [16], or on self-supervised objectives [17, 18, 19].
On the other hand, some methods [20, 21] train the CNN with temporal context by combining it with an LSTM network and training both components jointly on truncated video sequences. Training with temporal context rewards the CNN to extract features that could prove useful in hindsight and enables the CNN to tolerate ambiguous frames, relying on the LSTM to resolve these. Yet, these methods are limited to training on very short (few seconds) video sequences [20] or to CNNs without Batch Normalization (BN) [22] layers [21].
In contrast, we achieve to train a standard CNN with BN layers with extended temporal context (up to 64 frames) for visual feature extraction by using a particular, yet simple, strategy (see section 4.1): We make sure that the training batches are diverse by sampling several (at least ten) video sequences per batch from throughout the training data. After training, features are extracted using the CNN only.
1.0.2 Modeling long-range relationships
Attention mechanisms [8] were designed to model global relations in sequences. By stacking attention and feedforward neural networks, the Transformer [9] achieved break-through results on sequence modeling tasks. Inspired by this success, researchers incorporated attention in temporal models for phase recognition as well [23, 24, 25] (see Fig. 2, section 3.2).
However, specific challenges apply to the task of phase recognition, similar to the broader field of temporal action segmentation [26]: there are inherent constraints on locality and temporal order that need to be modeled, the videos are long and can vary widely in duration, and the available labeled datasets are small. Yet, attention mechanisms lack inductive bias and have quadratic time and memory complexity, which makes them computationally expensive on long sequences. Thus, using attention effectively and efficiently in temporal models for phase recognition is an area of ongoing research.
Previous methods chose strategies that constrain the capabilities of attention, namely special losses to control attention weights [23, 27] or reverting to local attention [24, 28, 27], which is computed within limited time windows and thus cannot capture global dependencies. In contrast, we propose a solution that incorporates global attention efficiently and without imposing hand-crafted constraints.
Therefore, we build upon the convolutional U-Net [29] encoder-decoder architecture, which was previously proposed for action segmentation [4, 30] but seemed recently less popular than multi-stage TCNs (MS-TCNs)[5]. A U-Net has the advantage of providing semantically meaningful features at the bottleneck after the encoder, which are also downsampled along the temporal dimension. We suggest to integrate attention at this position (see Fig. 2g, section 4.2), where attention can be computed efficiently and unconstrained. The proposed model, named TUNeS (Temporal U-Net with Self-Attention), is suited for both online and offline phase recognition.
1.0.3 Evaluation
We conducted a large number of experiments to demonstrate the positive effects of training the feature extractor with long temporal context (see Fig. 6). Our experiments also show that TUNeS compares favorably to many baseline methods, in terms of both recognition accuracy and computational efficiency (see Fig. 6 and Fig. 7). TUNeS, combined with a feature extractor that is trained with long temporal context, achieves state-of-the-art results on Cholec80 and AutoLaparo. The source code is available at https://gitlab.com/nct_tso_public/tunes.
2 Background
2.1 Surgical phase recognition
Given a video $v$ of length $T$, phase recognition is the task to estimate which phase $p$ is being executed at any time $t$. Here, $p$ is one of $C$ predefined surgical phases.
We denote the true phase that is happening at time $t$ as $y_{t}$, $1\leq y_{t}\leq C$, and the sequence of all ground truth labels as $\left(y_{t}\right)_{1\leq t\leq T}$ or $y$. The estimate of the phase at time $t$ is $\hat{y}_{t}\in\mathbb{R}^{C}$, where $\hat{y}_{t,p}$ is the score (i.e., the logarithm of the unnormalized probability) for the event that $y_{t}$ equals phase $p$. The video itself is a sequence of video frames $v_{t}$.
For intra-operative applications, phase recognition needs to be performed online. Here, only video frames $v_{t^{\prime}}$ up until the current time $t$, i.e., $t^{\prime}\leq t$, can be analyzed to estimate the current phase. For post-operative applications, however, offline recognition is feasible. Here, information from all frames in the video can be processed to estimate the phase at any time.
(a) Transformer block [9]
(b) Layer in MS-TCN [5] and ASFormer [28]
Figure 1: Related model blocks. ASFormer adds attention between the convolution and the fully connected layer of the MS-TCN layer.
Figure 2: Temporal models for surgical phase recognition or temporal action segmentation incorporate attention in different ways, see section 3.2. ($T$: length of visual feature sequence, $D$: dimension of visual feature, $C$: number of phases)
2.2 Attention mechanism
Let $Q=q_{1:S}$ be a sequence of queries and $U=u_{1:T}$ be a sequence of values, which is associated with a sequence $K=k_{1:T}$ of keys, i.e., $k_{t}$ is the key of $u_{t}$. In an attention mechanism [8], each element $q_{s}\in Q$ collects information from any element $u_{t}\in U$ by computing an attention-weighted sum over the complete sequence $U$:
$\begin{split}\mathrm{Attention}(Q,K,U)&=\left(x_{s}\right)_{1\leq s\leq S},\text{ where }\\ x_{s}&=\sum_{t}\frac{\exp{\alpha(q_{s},k_{t})}}{\sum_{t^{\prime}}\exp{\alpha(q_{s},k_{t^{\prime}})}}\cdot u_{t}\end{split}$ (1)
The weights in this sum are the attention weights $\alpha(q_{s},k_{t})$ after Softmax. Here, function $\alpha$ is trained to determine how relevant the element with key $k_{t}$ is given query $q_{s}$.
In practice, $Q$, $K$, and $U$ are passed through trainable linear layers before computing attention, and $\alpha$ simply computes the scaled dot product. In multi-head attention, multiple linear layers are used so that attention is computed on multiple representation subspaces in parallel. In the case of self-attention, $Q$, $K$, and $U$ all refer to the same sequence. [9]
Attention weights may be masked, i.e., set to $\infty$- ∞, according to specific rules: Local attention is only computed within local time windows of size $1+2\omega$, i.e., $\alpha(q_{s},k_{t})=-\infty$ for all $t$ with $|t-s|>\omega$. Causal attention does not allow attending to future sequence elements, thus $\alpha(q_{s},k_{t})=-\infty$ for all $t$ with $t>s$.
In the Transformer architecture [9], attention is embedded in Transformer blocks with residual connections and normalization layers (Fig. 1a). Here, the attention mechanism is followed by a pointwise non-linear feedforward network.
3 Related work
3.1 Phase recognition without attention mechanisms
Early models for video-based phase recognition consisted of a CNN for feature extraction and an LSTM network as temporal model [31, chapter 6.4]. Later, Czempiel et al. [14] proposed TeCNO for temporal modeling (Fig. 2a). TeCNO is a TCN with multiple stages [5], where subsequent stages refine predictions of previous stages. Each stage consists of several layers with dilated temporal convolutions (Fig. 1b), where the dilation factor doubles at each layer. For online recognition, causal dilated convolutions [32] are used, i.e., convolutions are computed over the current and preceding time steps only. Yi et al. [33] suggested to train multi-stage models not end-to-end, but to apply disturbances to the intermediate predictions during training. Recently, Kadkhodamohammadi et al. [34] introduced PATG, which uses a SE-ResNet as feature extractor and a graph neural network as temporal model. Zhang et al. [35] proposed a reinforcement learning approach to retrieve phase transitions for offline recognition.
Whereas feature extractor and temporal model were trained separately in these studies, others proposed to train both models jointly. Jin et al. [20] trained SV-RCNet, a ResNet-LSTM, end-to-end on 2 s long video sequences. Later, they proposed TMRNet [36], which augments a ResNeSt-LSTM with a long-range memory bank [37] so that feature vectors from the previous 30 s can be queried in addition. Recently, Rivoir et al. [21] showed that CNN-LSTMs should be trained on even longer video sequences, which becomes feasible only if effects caused by BN are taken into account. They trained their best model, a ConvNeXt-LSTM, on sequences of 256 s.
3.2 Phase recognition with attention mechanisms
To model long-range dependencies, recent studies proposed to integrate attention mechanisms into temporal models (Fig. 2). With attention, an element $q_{s}$ in the query sequence can be related to any element $u_{t}$ in the value sequence, regardless of their temporal distance $|t-s|$.
Czempiel et al. [23] presented OperA, a stack of Transformer blocks, for online recognition (Fig. 2b). This approach uses only self-attention to model temporal dependencies and thus may be lacking a notion of locality and temporal order. Since the model seemed to struggle to filter relevant information from the feature sequences, the authors proposed to train with a special attention regularization loss.
Gao et al. [24] introduced Trans-SVNet for online recognition (Fig. 2c), which adds two Transformer blocks on top of a frozen TeCNO model. The first block computes self-attention on the TeCNO predictions. In the second block, the initial predictions of the feature extractor attend over the output of the first Transformer block. All attention operations are computed locally, using the feature vectors of the previous 30 s only.
Yi et al. [28] proposed ASFormer, an offline model for action segmentation (Fig. 2d). This model integrates attention into MS-TCN layers (Fig. 1b), where the convolutions help to encode local inductive bias. Again, attention is computed within local windows. Here, the window size equals the dilation factor $d$ of the convolution, which doubles at each layer. Restricting attention to local neighborhoods of a predefined size may help to handle the complexity of long, noisy feature sequences but hinders the modeling of global dependencies.
Recently, Zhang et al. [38] suggested to add a MS-TCN stage in parallel to the first ASFormer stage and to fuse their outputs. Chen et al. [39] presented an online two-stage version of ASFormer, where the dilation factors in the second stage halve (instead of double) at each layer.
Whereas previous models incorporated hierarchical patterns implicitly, namely by means of doubling dilation factors, Ding and Li [25] proposed an explicitly hierarchical model: SAHC (Fig. 2e). This model uses MS-TCN-like stages to process feature sequences, which are downsampled progressively. Then, the obtained multi-scale feature maps are fused from coarse to fine. Transformer blocks with global attention are integrated at the very end, where the final high-resolution sequence queries further information from the lower-resolution sequences.
Du et al. [27] presented Temporal U-Transformer (TUT), a hierarchical offline model for action segmentation (Fig. 2f). TUT consists of Transformer blocks that are equipped with downsampling and upsampling layers to form a U-Net. Here, blocks in the upsampling path obtain information from the downsampling path by attending over the respective feature maps. Multiple U-Nets are stacked to create a multi-stage architecture. Local attention is used throughout the model. In addition, a boundary-awareness loss is employed to control the distribution of attention weights around action boundaries.
4 Methods
4.1 Feature extractor with temporal context
We use a standard CNN as feature extractor: ResNet-50 [10], pre-trained on ImageNet. However, we train the CNN not on individual video frames (Fig. 3a) but with temporal context. For this purpose, we add an LSTM cell with 512 hidden units, which processes the sequence of computed CNN features. The CNN-LSTM is trained end-to-end on sequences of $L$ consecutive frames (Fig. 3b). Here, we minimize the cross-entropy loss, computed on all $L$ frames in the sequence.
To train on a balanced set of video sequences, we sample five sequences from each phase and one sequence around each phase transition from each video during one training epoch. To avoid BN-related problems, video sequences are processed in appropriately large batches of $B\times L$ frames, $B\geq 10$.
After training, only the CNN is used to extract features: For each frame $v_{t}$, we obtain the feature vector $f_{t}\in\mathbb{R}^{D}$, $D=2048$, after the ResNet’s global average pooling layer. To model global dependencies and for offline recognition, a temporal model is trained on the extracted feature sequences $\left(f_{t}\right)_{1\leq t\leq T}$ in a second step.
Figure 3: Feature extractor training: (a) Training the CNN on individual video frames. The fully connected (FC) layer acts as linear classifier on top of the extracted feature vector $f_{t}$. (b) Training the CNN with temporal context, using an LSTM cell to model temporal relationships.
4.2 Temporal U-Net with self-attention
For temporal modeling, we propose to combine global self-attention to model long-range dependencies and a temporal convolutional U-Net to introduce local inductive bias.
By design, a U-Net extracts a hierarchy of multi-scale feature maps. We position Transformer blocks at the bottleneck after the U-Net encoder to compute self-attention on the highest-level features, which presumably are pre-filtered and semantically meaningful and thus adequate for attention mechanisms. Here, the feature maps are also temporally downsampled by an order of magnitude so that the computational cost of attention is acceptable. In addition, global dependencies are analyzed at the core of the U-Net and can be refined further in the upsampling path. This is different in Trans-SVNet and SAHC, which integrate attention towards the end of the model. Notably, we do not constrain the flexibility and expressivity of attention by using regularization losses (like OperA, TUT) or local attention only (like Trans-SVNet, ASFormer, TUT).
The proposed model is a temporal U-Net with self-attention: TUNeS (Fig. 4). Previously, similar architectures were used successfully for image generation [40, 41], image segmentation [42, 43], or speech denoising [44].
Figure 4: TUNeS is a U-Net of temporal convolutions, which consists of seven TUNeS blocks. To model long-range dependencies, self-attention is used at the bottleneck (block 4). TUNeS processes the sequence of visual features $f$, which is obtained by applying the feature extractor (section 4.1) to every frame $v_{t}$ (with dimensions $H\times W$) in the video. Here, the video is depicted as videogram [45]. Multi-scale phase predictions $\hat{y}^{(i)}$ are extracted from the U-Net decoder to supervise the model during training. ($T$: length of video, $D$: dimension of visual feature, $C$: number of phases)
4.2.1 Model architecture
In the encoder, the feature sequence $\left(f_{t}\right)_{1\leq t\leq T}$ is repeatedly passed through temporal convolutions and downsampled. In the decoder, the sequence is repeatedly upsampled and passed through further convolutions. At the bottleneck, TUNeS Transformer blocks (Fig. 5a) with multi-head self-attention enrich the sequence with global context. Here, the downsampled feature sequence is treated as a sequence of tokens, to which start and end tokens are added. The TUNeS Transformer block starts with a convolutional block to add conditional positional encodings [46].
In the decoder, the intermediate feature sequence is enriched with the information of higher-resolution feature maps by means of skip connections to the encoder. Here, the sum in each skip connection is weighted with a learnable scalar. Further, linear classifier heads after each decoder block compute phase predictions at multiple temporal scales.
Encoder and decoder consist of convolutional blocks, which are MS-TCN layers (Fig. 1b) with GELU [47] activation. Down- und upsampling by factor f is implemented by 1D convolutions with kernel f and stride f, where the convolution is transposed for upsampling. Dilated convolutions ($d=18$) are used in the decoder to retain a large receptive field when progressing to higher-resolution sequences. We set $d=18$ since the coarsest feature maps are downsampled by this factor.
TUNeS does not include any normalization layers. Each operation, including each attention head, processes and outputs sequences with $\mathrm{dim}=64$ channels.
4.2.2 Causal operations for online recognition
To be applicable to online recognition, it is required that the temporal model does not access future information. Thus, TUNeS uses causal convolutions for online recognition like TeCNO [14]. In addition, causal self-attention is computed in the Transformer blocks by applying a causal mask to the attention weights.
It is important to perform the downsampling operations in a causal way as well. Like for convolutions, this can be implemented by shifting input sequences by $k-1$ elements to the right, which corresponds to adding $k-1$ padded elements at the left (Fig. 5b). Here, $k$ is the size of the kernel. In contrast, regular downsampling would leak future information. In that case, an element in the output sequence could depend on up to $k-1$ future elements in the input sequence.
4.2.3 Alternate attention masking for offline recognition
When computing global self-attention in offline mode, it may be difficult to correctly separate past from future information. To help with this, we alternately prevent access to future and to past elements in the TUNeS Transformer blocks (Fig. 5c). Further, we found it beneficial to increase the number of Transformer blocks for offline recognition from two to eight.
(b) Regular vs. causal operations
(c) Alternate attention masking
[…]
4.2.4 Training objective
Let $\left\{\hat{y}^{(i)}:1\leq i\leq 4\right\}$ be the multi-scale phase predictions, where the length of prediction $\hat{y}^{(i)}$ is reduced by factor $\chi_{i}$ and $\chi=(1,3,9,18)$. For $2\leq i\leq 4$, we compute downsampled annotations $y^{(i)}=(y^{(i)}_{s})_{1\leq s\leq T/\chi_{i}}$, where $y^{(i)}_{s}\in\{0,1\}^{C}$ is a binary vector that summarizes the section $\{y_{t}:(s-1)\chi_{i}<t\leq s\chi_{i}\}$ of $\chi_{i}$ frames in the full-resolution annotation: If phase $p$ appears in the section, the $p$-th entry in $y^{(i)}_{s}$ is 1 and 0 otherwise. Since a section may cover more than one phase, $y^{(i)}_{s}$ may have more than one entry that is non-zero. Thus, the downsampled annotations are multi-label annotations.
The training objective $\mathcal{L}_{\text{total}}$ summarizes losses at all scales, see (2). We use cross-entropy (CE) loss for the full-resolution prediction $\hat{y}^{(1)}$. For the lower-resolution predictions, we use binary cross-entropy (BCE) loss to handle the multi-label annotations. To counter class imbalance, we use class weights $\gamma$, which are computed on the training data following the median frequency balancing [48] approach. In addition, we compute a smoothing loss $\mathcal{L}_{\text{smooth}}$ [5] on the full-resolution prediction, which is weighted by factor $\lambda_{\text{smooth}}=0.15$.
$\begin{split}\mathcal{L}_{\text{total}}=\,\,&\mathcal{L}_{\text{phase}}(\hat{y}^{(1)},y)+\sum_{k=2}^{4}\mathcal{L}_{\text{BCE}}(\hat{y}^{(k)},y^{(k)})\\ &\mathcal{L}_{\text{phase}}(\hat{y},y)=\mathcal{L}_{\text{CE}}(\hat{y},y)+\lambda_{\text{smooth}}\cdot\mathcal{L}_{\text{smooth}}(\hat{y})\end{split}$ (2)
The individual loss terms are defined in the supplementary.
4.2.5 Regularization
For data augmentation, we augment feature sequences by (i) shifting the complete sequence to the right by a small random number of time steps and by (ii) increasing and decreasing the video speed by randomly dropping and duplicating elements of the feature sequence.
In addition, we mask spans in the token sequences that are input to the Transformer blocks, inspired by the pre-training procedures for masked language models [49, 50]. Here, we randomly sample masks to cover spans that correspond to 18–300 s in the full-resolution video. However, we make sure that phase transitions are not masked because it would be impossible to reconstruct their exact timing. In total, around 35% of all tokens are masked per sequence.
5 Experiments
5.1 Dataset and evaluation metrics
Experiments were conducted on the Cholec80111http://camma.u-strasbg.fr/datasets dataset [13]. Cholec80 consists of 80 video recordings of laparoscopic gallbladder removals, which are labeled with phase and tool information. Here, $C=7$ different surgical phases are annotated. We used only the phase labels to train feature extractors and temporal models. The videos were processed at a temporal resolution of 1 fps. Unless stated otherwise, experiments were performed on the 32:8:40 data split, meaning that the first 32 videos in Cholec80 were used for training, the next 8 videos for validation, and the last 40 videos for testing. In particular, hyperparameters were tuned on the validation subset.
We use the following evaluation metrics: Video-wise accuracy, video-wise Macro Jaccard (or Intersection over Union), and $\mathit{F1}$ score, which is computed as the harmonic mean of the mean video-wise Precision and the mean video-wise Recall. The Macro Jaccard for one test video is the average of the phase-wise Jaccard scores computed on this video. To avoid problems with undefined values, we omit the phase-wise scores for phase $p$ and video $v$ if $p$ is not annotated in $v$.
For comparison with prior art, we also report the relaxed evaluation metrics -$\mathcal{R}\text{-}\mathsf{Jaccard}$ (or -$\mathcal{R}\text{-}\mathsf{Precision}$ and -$\mathcal{R}\text{-}\mathsf{Recall}$) and -$\mathcal{R}\text{-}\mathsf{Accuracy}$, which were proposed to tolerate certain errors within 10 s before or after an annotated phase transition. For all metrics, we used the implementation provided by [51].
5.2 Comparisons to baselines and state of the art
We conducted experiments 1) to explore the impact of providing longer temporal context when training the visual feature extractor, and 2) to compare TUNeS to an extensive number of established temporal models.
5.2.1 Feature extractors
We trained feature extractors with context lengths $L$ of 1, 8, or 64 frames (Fig. 6). The feature extractor with $L=1$ is simply a ResNet-50, without the LSTM on top (Fig. 3a). Following the custom sampling strategy (section 4.1), 1264 video sequences were processed in each epoch, for any size of $L$. With $L=64$, we trained the feature extractor for 200 epochs. For comparability, we trained the feature extractor with $L=8$ for 1600 epochs and with $L=1$ for 12800 epochs. Thus, all feature extractors see the same number of video frames during training, but in different contexts and in different order. The batch size $B$ was 10 video sequences with $L=64$, 80 with $L=8$, and 128 video frames with $L=1$.
To assess the feature extractor’s abilities when $L>1$, we used the predictions of the jointly trained LSTM: To infer the current phase, we applied the CNN-LSTM to the $L-1$ previous and to the current frame and took the final prediction. In addition, we performed carry-hidden evaluation (CHE) [21], where the LSTM’s hidden state is carried through the video.
To account for the randomness in training deep-learning models, experiments were repeated five times using different random seeds.
5.2.2 Temporal models
For each context length $L\in\{1,8,64\}$, we trained TUNeS as well as other temporal models on the extracted feature sequences (Fig. 6). The baseline models included an RNN (2-layer GRU), an MS-TCN (TeCNO), and models that integrate attention in different ways. For comparability, we set the hidden dimension of all models to $\mathrm{dim}=64$. To account for randomness, we repeated each experiment five times on each of the five feature extractor instances, yielding 25 experimental runs in total.
All models were trained for 75 epochs using the Adam [52] optimizer and a batch size of one feature sequence. Baseline models were trained on the combined loss $\mathcal{L}_{\text{phase}}$ with $\lambda_{\text{smooth}}=0.15$, see (2). For models with multiple stages, we computed the loss on the output of all stages. Feature sequence augmentation was applied as described in section 4.2.5. Gradients were clipped to have a maximum norm of 1. Finally, we selected the model that achieved the highest Macro Jaccard on the validation set for testing.
To select an appropriate learning rate $\eta$ for each model, we ran a quick sweep over $\eta\in\{10^{-4},3\cdot 10^{-4},5\cdot 10^{-4},10^{-3}\}$. Usually, $\eta=5\cdot 10^{-4}$ worked well. If available, we tried to consider the training strategy of the original code base. Full training details are listed in the supplementary material.
[…]
Figure 6: Surgical phase recognition results on the Cholec80 dataset (32:8:40 split). We present the feature extractor’s standalone performance as well as the performance of various temporal models, trained on top of the frozen feature extractor. Error bars indicate the standard deviation over repeated experimental runs. The asterisk* indicates training without feature sequence augmentation and smoothing loss. Numerical results can be found in the supplementary material.
5.2.3 Performance measurements
For TUNeS as well as for the baselines, we measured model latency and peak GPU memory allocation in offline mode. To this end, we used synthetic feature sequences, which are randomly initialized tensors of dimension $T\times 2048$, with a length $T$ of 450, 900, 1800, 3600, and 7200 frames. Fig. 7 presents the mean of 1000 measurements, after taking 100 measurements for warm-up.
5.2.4 Comparison to the state of the art
We compare the results of TUNeS, trained on features with maximum context $L=64$, with previously reported results on the 40:40 split (Tables 1 and 2). Here, the feature extractor with $L=64$ and TUNeS were re-trained on all of the first 40 videos in Cholec80. Due to the lack of a validation set, we selected the model after the last epoch for testing. To improve convergence, we continuously reduced the learning rate during training, following a cosine annealing schedule [53]. For comparisons on the 40:20:20 split, we evaluated the models trained for the 40:40 split on the last 20 videos in Cholec80.
[…]
Figure 7: Performance measurements for processing feature sequences of increasing duration with different temporal models in offline mode. We also present the number of parameters for each temporal model.
Discussion
Training the feature extractor in context proved beneficial: The standalone performance of the feature extractor improved as the temporal context increased. In addition, almost all temporal models performed better on top of feature extractors that were trained with longer temporal context (Fig. 6). The only exception is TUT, which performed better on features with $L=8$ than on features with $L=64$.
Apparently, training in context helps the CNN to learn meaningful, contextualized features that hold crucial information for subsequent analysis. Yet, training in context is simple to implement and does not require additional manual annotations such as tool labels. Also, it does not take more time than the standard frame-wise training. The only requirement is sufficient GPU memory (32 GB in our experiments).
When evaluating the baselines (Fig. 6), we found that the 2-layer GRU and OperA achieved subpar results. For the GRU, we suspect that regularization with the smoothing loss and the proposed feature sequence augmentation is not as effective for a recurrent model. For OperA, we presume that the model is suffering from a lack of local inductive bias and struggles to filter information from noisy features. Yet, OperA improves considerably on contextualized features ($L=64$), which might be more meaningful and thus easier to analyze with attention. Please note that we used a custom implementation of OperA because official code has not been released.
TUT uses similar building blocks as OperA but performed considerably better by processing multi-scale sequences with local attention. Yet, TUT did not reach the performance of ASFormer or TUNeS. Since TUT downsamples and upsamples the feature sequence but abandons skip connections in favor of cross-attention, it may lose too much detail information.
Training Trans-SVNet on top of TeCNO improved the results only marginally. Yet, we tuned hyperparameters for TeCNO, e.g., the number of layers per stage, and used a larger model dimension and more training epochs than in the original implementation. We assume that Trans-SVNet mostly learns to locally smooth the TeCNO outputs, which may not be necessary if TeCNO was tuned and trained properly.
The best performing baselines were SAHC, TeCNO, TUT, and ASFormer. TUNeS outperformed the baselines in terms of accuracy and scored better or similarly on Macro Jaccard (Fig. 6). The results achieved by TUNeS, trained on top of the feature extractor with $L=64$, also compare favorably to previous results from the literature (Tables 1 and 2). Moreover, TUNeS is computationally efficient and scales well on long sequences of up to two hours, regarding both latency and memory consumption (Fig. 7).
Table 1: Online recognition on Cholec80 (40:40 split). Comparison to results from the literature.
Model -$\mathcal{R}\text{-}\mathsf{Accuracy}$$\mathcal{R}\text{-}\mathsf{Jaccard}$
$M$ $SD_{V}$ $M_{P}$ $SD_{P}$
ResNet-50–LSTMb, $L=8$ (similar to SV-RCNet[20]) 0.881 0.063 0.739 0.090
TeCNO [14], results of [24] 0.886 0.078 0.751 0.069
TMRNet w/ ResNeST [36] 0.901 0.076 0.791 0.057
Trans-SVNet [24] 0.903 0.071 0.793 0.066
Not E2E [33] 0.920 0.053 0.771 0.115
SAHC [25] 0.918 0.081 0.812 0.055
ConvNeXt–LSTMb, $L=256$ [21] 0.935 0.065 0.829 0.101
TUNeS, $L=64$ 0.939 0.050 0.842 0.083
Like prior art, we report the mean $M$ and the standard deviation over videos $SD_{V}$ for -$\mathcal{R}\text{-}\mathsf{Accuracy}$. For -$\mathcal{R}\text{-}\mathsf{Jaccard}$, we report the mean over phase-wise means $M_{P}$ and the standard deviation over phases $SD_{P}$.
Model Accuracya Macro Jaccarda $\mathit{F1}$ scorea
ResNet-50–LSTMb, $L=8$ 0.872 $\pm$ 0.004 0.707 $\pm$ 0.007 0.832 $\pm$ 0.003
TeCNO [14]
results of [11] 0.874 $\pm$ 0.014 0.825 $\pm$ 0.018
results of [39] 0.900 0.723 0.839
Trans-SVNet [24]
results of [39] 0.896 0.731 0.845
GRU w/ VideoSwin [12] 0.909 0.853
PATG w/ SE- ResNet-50 [34] 0.914 0.854
Dual Pyramid ASFormer w/ PCPVT [39] 0.914 0.754 0.858
TUNeS, $L=64$ 0.927 $\pm$ 0.004 0.799 $\pm$ 0.005 0.890 $\pm$ 0.003
aThe reported sample standard deviation refers to the variation over repeated experimental runs.
bTrained end-to-end on short video sequences of $L$ frames.
Table 2: Offline recognition on Cholec80. Comparison to results from the literature.
Model -$\mathcal{R}\text{-}\mathsf{Accuracy}$$\mathcal{R}\text{-}\mathsf{Precision}$$\mathcal{R}\text{-}\mathsf{Recall}$ split $M$ $SD_{V}$ $M_{P}$ $SD_{P}$ $M_{P}$ $SD_{P}$ ASTCFormer [38] 0.957 0.033 0.923 0.062 0.912 0.095 40:20:20 TUNeS, $L=64$ 0.964 0.022 0.936 0.070 0.939 0.054 40:20:20 Like prior art, we report the mean $M$ and the standard deviation over videos $SD_{V}$ for -$\mathcal{R}\text{-}\mathsf{Accuracy}$. For -$\mathcal{R}\text{-}\mathsf{Precision}$ and -$\mathcal{R}\text{-}\mathsf{Recall}$, we report the mean over phase-wise means $M_{P}$ and the standard deviation over phases $SD_{P}$.
Model Accuracya Macro Jaccarda $\mathit{F1}$ scorea split GRU w/ VideoSwin [12] 0.939 0.898 40:40 TUNeS, $L=64$ 0.949 $\pm$ 0.005 0.836 $\pm$ 0.009 0.912 $\pm$ 0.005 40:40 Transition Retrieval Network [35] 0.901 0.848 40:20:20 TUNeS, $L=64$ 0.953 $\pm$ 0.004 0.837 $\pm$ 0.011 0.911 $\pm$ 0.007 40:20:20 aThe reported sample standard deviation refers to the variation over repeated experimental runs.
TUNeS is computationally lightweight due to its U-Net structure, where attention is computed only at the bottleneck. In particular, TUNeS is faster than TeCNO, TUT, and SAHC. OperA has even lower latency, but struggles with increased memory usage on long sequences due to the quadratic memory complexity of attention. In contrast, ASFormer uses a custom implementation of local attention to restrict memory consumption, which comes at the cost of increased computation times, especially on long sequences. Due to the sequential computation, GRU is also relatively slow on long sequences.
[…]
Figure 8: Impact of integrating Transformer blocks at different positions in the TUNeS model. Left: Results on the Cholec80 dataset (32:8:40 split). Error bars indicate the standard deviation over repeated experimental runs. Right: Performance measurements in offline mode.
OperAw/o attention maskingSAHCw/ duplicated Transformer blocksw/ duplicated Transformer blocks + attention maskingTUNeSw/o attention masking1864Features’ context $L$ (s).80.85.90.95AccuracyOffline recognition1864Features’ context $L$ (s).60.70.80Macro Jaccard
Figure 9: Impact of attention masking so that Transformer blocks alternately process past and future information in offline mode (Fig. 5c). We present results on the Cholec80 dataset (32:8:40 split). Error bars indicate the standard deviation over repeated experimental runs.
5.3 Ablation studies
For further insights into TUNeS, additional experiments were conducted on the 32:8:40 split. In each ablation, one specific aspect was varied as described below.
5.3.1 Impact of the position of Transformer blocks
Originally, Transformer blocks are incorporated at TUNeS block 4, see Fig. 4. We studied positioning Transformer blocks at other TUNeS blocks $b$ before ($b\in\{1,2,3\}$) or after ($b\in\{5,6,7\}$) the bottleneck. Here, we replaced the convolutional blocks in block $b$ with $n$ Transformer blocks (online mode: $n=2$, offline mode: $n=8$) and used two convolutional blocks at the bottleneck. Thus, we swapped TUNeS blocks $b$ and 4, meaning that the overall number of model parameters remained unchanged. As lower bound, we trained a variant of TUNeS that does not use attention at all. Here, we replaced the $n$ Transformer blocks at the bottleneck with $n$ convolutional blocks. For all model configurations, we conducted performance measurements as described in section 5.2.3.
[…]
Figure 10: A 2-stage variant of TUNeS can be used to reduce over-segmentation errors. Left: Offline recognition results on the Cholec80 dataset (32:8:40 split). Error bars indicate the standard deviation over repeated experimental runs. Right: Performance measurements in offline mode.
Discussion
We hypothesize that Transformer blocks should be positioned at the bottleneck so that attention (i) is computed on high-level features, and (ii) is used early so that global context is available throughout the decoder. The results (Fig. 8), support our intuition: If Transformer blocks are positioned further away from the bottleneck, recognition accuracy deteriorates. In offline mode, where more Transformer blocks are used, training becomes more unstable. Also, attention needs to be computed on higher-resolution sequences, which is more costly regarding latency and GPU memory.
Notably, using Transformer blocks towards the end of the model, where features are expected to be more meaningful, performs better than using Transformer blocks at the beginning. Yet, on highly contextualized features ($L=64$) this effect is less pronounced in online mode. Compared to TUNeS, the model with no attention at all performs considerably worse. However, using no attention is often better than using attention at ineffective positions, especially in TUNeS blocks 1 and 2.
5.3.2 Impact of alternate attention masking
We trained and tested models for offline recognition either with attention masks, which alternately prevent access to future and to past information, or without any attention masks. Here, we studied three temporal models with global attention: OperA, SAHC, and TUNeS. For SAHC, which uses individual Transformer blocks instead of stacks of multiple Transformer blocks (Fig. 2e), we duplicated each Transformer block to implement alternate attention masking. Then, the first block queries past information and the second block queries future information.
Discussion
Even though attention masks restrict the information that can be accessed in each Transformer block, all temporal models achieved better results with attention masks, see Fig. 9. This applied also to the advanced SAHC model. Here, an additional experiment with duplicated Transformer blocks but without attention masks demonstrates that the improvement does not come from additional model parameters alone. For OperA, which is lacking local inductive bias, alternate attention masking is actually crucial for acceptable performance. Notably, He et al. [12] also reported difficulties to train a 1-layer Transformer encoder, which is similar to the OperA model, without attention masks for offline recognition.
5.3.3 Quality of predicted phase segments
The frame-wise annotation $y$ induces a partitioning $Y$ into phase segments, where each segment summarizes all consecutive frames of a phase. Likewise, the frame-wise prediction $\hat{y}$ induces the estimated segmentation $\tilde{Y}$. Due to the incorrect classification of individual frames, it is common for $\tilde{Y}$ to divide a video into many fragmented segments. To quantify this, Lea et al. [54] proposed the edit score, which is the normalized Levenshtein distance between $Y$ and $\tilde{Y}$. This metric penalizes out-of-order predictions and over-segmentation errors.
Fig. 10 shows the mean video-wise edit score for phase segments predicted in offline mode by TUNeS and other temporal models. TUNeS obtains higher edit scores than GRU, OperA, and SAHC. Yet, the 2-stage models yield better segmentations, especially TUT. We assume that the multi-stage setup – in addition to the repeated down- and upsampling in TUT – helps to reduce noise and segment fragmentation. Consequently, we also trained a 2-stage variant of TUNeS, see (3). Here, we optimized the sum of the losses at both stages.
[…]
2-stage TUNeS achieves high edit scores that approach those of TUT and, in contrast to TUT, preserves high precision in terms of accuracy and Macro Jaccard. However, this improvement comes at the cost of increased computation time (Fig. 10, right), which doubles due to the additional TUNeS stage.
In online mode, where the current phase needs to be recognized immediately, with incomplete information, and incorrect decisions cannot be revised, it is difficult to avoid noisy predictions. Here, initial experiments showed that 2-stage TUNeS could also improve the edit score, but this could come at the cost of decreased accuracy and Macro Jaccard. In addition, even the best models with regard to edit score (GRU and 2-stage TUNeS) could not reach a score above $40\,\%$.
Online recognitionOffline recognition.91.93.95TUNeS (as proposed)No convolution inTransformer blocksSinusoidalpositional encodingNo dilated convolutionsin decoder ($d=1$)No dropoutNo token maskingNo feature sequenceaugmentationFeatures extractedfrom LSTM (not CNN)No attention(convolutions only).923.917.918.919.921.919.917.912.906.946.945.931.945.945.940.945.932.916Accuracy.77.80.83.784.780.779.778.783.778.770.772.765.825.827.804.826.823.823.816.803.791Macro Jaccard
Figure 11: Results of ablation studies on the Cholec80 dataset (32:8:40 split). In all studies, we used features with context $L=64$. Error bars indicate the standard deviation over repeated experimental runs.
5.3.4 Additional ablations
For $L=64$, we conducted ablation experiments with the following modifications (Fig. 11 and 12): (i) To explore the importance of conditional positional encodings, we (a) omitted the convolutional block in the TUNeS Transformer block or (b) used standard sinusoidal positional encodings [9] instead. (ii) We used regular instead of dilated convolutions in the TUNeS decoder. (iii) To investigate the effect of measures for regularization, we (a) trained without dropout, i.e., we deactivated dropout in the convolutional blocks and feedforward networks, (b) trained without token masking, or (c) trained without feature sequence augmentation. (iv) We extracted the visual features from the LSTM cell of the feature extractor, using the features right before the final linear classifier. (v) To study the importance of attention, we replaced the TUNeS Transformer blocks with convolutional blocks. Additional results for TUNeS without attention are presented in Fig. 8. (vi) To explore the impact of the number $n$ of Transformer blocks, we varied $n\in\{2,4,6,8,10\}$.
Discussion
As shown in Fig. 11, most modifications impaired the results, but in many cases the decline was not severe.
(i) Omitting the convolutional positional encoding did not lead to considerably worse results. Probably, the convolutions in the encoder provide sufficient local inductive bias. Also, sinusoidal positional encodings did not improve results. In online mode, they worked as well as using no positional encodings. In offline mode, they caused training instability and overall inferior results. Presumably, this kind of rigid and absolute positional encoding is not suitable for modeling past and future information in sequences of widely varying length. (ii) Foregoing dilations in the decoder convolutions slightly affected the results for online recognition. The offline model seems to be robust in this regard, possibly due to the acausal operations, which have a symmetric receptive field. (iii) Regarding different regularization techniques, we found that deactivating dropout had only a minimal effect on online and offline recognition. Dropping token masks worsened the results slightly. Omitting feature sequence augmentation had a noticeable negative effect. In particular, Macro Jaccard dropped by 1.8 % for online recognition and by 1.1 % for offline recognition. (iv) Interestingly, training TUNeS on LSTM features instead of CNN features impaired the results considerably. Possibly, some of the detail in the CNN features is lost during additional processing with the LSTM. Further, the LSTM features might fit the training data too well, making it difficult for the temporal model to learn generalizable patterns based on that. (v) The impact of attention was the greatest: Without attention to model global temporal relationships, Macro Jaccard dropped by 2.4 % in online mode and by 4.1 % in offline mode. Without attention, TUNeS performs similarly to TeCNO, which might indicate that multi-scale modeling alone cannot explain the performance of TUNeS. Rather, the effective integration of Transformer blocks is a crucial factor. (vi) When scaling the number of Transformer blocks (Fig. 12), we found that using more Transformer blocks was helpful for offline recognition but not for online recognition. Generally, it might be difficult to further improve on the online task due to the inherent uncertainty and ambiguity of this problem. Here, even global information is incomplete because it is only information from the past.
.91.93.95Accuracy.923.924.924.925.924Online recognition.929.939.942.946.942Offline recognition246810Number ofTransformer blocks.77.80.83Macro Jaccard.784.782.779.776.773246810Number ofTransformer blocks.800.812.821.825.820
Figure 12: Impact of the number of Transformer blocks used in TUNeS. We present results on the Cholec80 dataset (32:8:40 split), using features with context $L=64$. Error bars indicate the standard deviation over repeated experimental runs.
To conclude, the ablation studies helped us identify three major aspects that are decisive for the performance of TUNeS:
  • enriching the temporal U-Net structure with attention to model long-range relationships,
  • integrating attention at an appropriate position in the U-Net structure, and
  • using alternate attention masking for offline recognition to help separate past and future information.
[…]
Figure 13: Results on the AutoLaparo dataset (10:4:7 split). We present the performance of the feature extractor (ResNet-50–LSTM) and of TUNeS, trained on features with context $L=64$. Error bars indicate the standard deviation over repeated experimental runs.
5.4 Evaluation on the AutoLaparo dataset
We additionally evaluated TUNeS on AutoLaparo [^2] https://autolaparo.github.io [55], which consists of 21 video recordings of laparoscopic hysterectomies. This dataset provides manual annotations of $C=7$ surgical phases at a resolution of 1 fps. Following the proposed data split, we used the first 10 videos for training, the following 4 videos for validation, and the last 7 videos for testing.
We followed section 5.2.1 to train feature extractors with context lengths $L\in\{8,64\}$. Here, we doubled the number of training epochs to account for the smaller number of training videos and thus training sequences that would be sampled per epoch. On the features with $L=64$, we trained TUNeS exactly as described in section 5.2.2, but using 100 training epochs. Here, we trained with learning rate scheduler and selected the model after the last epoch for testing. In addition, we re-trained TUNeS on the combined set of training and validation videos.
The results are presented in Fig. 13. To compare to [55], we additionally report $f\text{-}\mathsf{Jaccard}=1/C\sum_{p=1}^{C}f\text{-}\mathsf{Jaccard}_{p}$. Here, the true positives, false positives, and false negatives to compute the phase-wise $f\text{-}\mathsf{Jaccard}_{p}$ are counted over all frames in the test set instead of individually for each video.
Training the feature extractor with long temporal context yielded major improvements over the baselines. TUNeS, trained on top of the contextualized features, achieved additional improvements over the standalone ResNet-LSTM. Here, we obtained a mean video-wise accuracy of 0.851 and a mean video-wise Macro Jaccard of 0.685 for online recognition. Yet, we see potential for improvement on the hysterectomy phase recognition task once more training data becomes available. For a start, TUNeS clearly benefits from including the four videos from the validation set during training.
6 Conclusion
In this paper, we proposed a simple method for training a standard CNN with longer temporal context to extract more meaningful visual features for surgical phase recognition at no added cost (section 4.1). We showcased the superior performance of feature extractors with extended context on the Cholec80 and AutoLaparo datasets.
Moreover, we presented TUNeS (section 4.2), a computationally efficient temporal model for surgical phase recognition, which achieves competitive results on Cholec80 in both online and offline mode.
TUNeS combines convolutions at multiple temporal resolutions, which are arranged in a U-Net structure, with Transformer blocks to model long-range temporal relationships. Importantly, the Transformer blocks are integrated at the core of the U-Net, where they can be employed most effectively and efficiently. In addition, the convolutions encode local inductive bias and the proposed alternate attention masking supports a notion of temporal order in offline mode. Therefore, TUNeS does require neither hand-crafted local attention patterns nor custom attention regularization losses – in contrast to previous temporal models that integrated Transformer blocks. These results contribute to a deeper understanding of how to leverage attention mechanisms for surgical phase recognition.